a study conducted at a certain college shows that 65% of the school's graduates find a job in their chosen field within a year of graduation. find the probability that among 5 randomly selected graduates, at least one finds a job in his or her chosen field within a year of graduating.
Answer by stanbon(75887)

a study conducted at a certain college shows that 65% of the school's graduates find a job in their chosen field within a year of graduation. find the probability that among 5 randomly selected graduates, at least one finds a job in his or her chosen field within a year of graduating.
-------------------------------------
Binomial Problem with n = 5 and p = 0.65
---
P(at least one finds job) = 1 - P(none find job)
---
= 1 - 0.35^5
---
= 1 - 0.0053
---------
= 0.9947
=============
Cheers,
stan H.